OpenStudy (anonymous):

State whether the relationship between the variables in the table is a direct variation, an inverse variation, or neither. If it is a direct or inverse variation, write a function to model it. x=3,9,13,20 y=9,27,39,60 I know it's direct but how do I write a function model to it?

hartnn (hartnn):

check : if the product xy is constant for all pairs of x,y then the variation is inverse, but if the ratio y/x is constant for all pairs of x,y then the variation is direct.

OpenStudy (anonymous):

but I don't get how to write a function model to it? :/

hartnn (hartnn):

so did u get it as inverse or direct ?

OpenStudy (anonymous):

direct? it says in my notebook if both x and y increases then it's direct

OpenStudy (anonymous):

ooh, do i just put the numbers into the equation y=k/x?

hartnn (hartnn):

yes, here if you note then, y/x = 9/3 = 3 y/x = 27/9 =3 y/x = 39/13 =3 so , you see everytime the ratio y/x turns out to be = 3 hence your function will also be y/x =3 or y =3x.

hartnn (hartnn):

yes, thats what i did, y= kx for direct . 9=k *3 ---> k=3 so, y=3x :)

OpenStudy (anonymous):

I get it now! thank you very much :D xx

hartnn (hartnn):

welcome ^_^